3 x + 5 y = 2
9 x - 5 y = 9

Answer by Alan3354(69443)

3 x + 5 y = 2
9 x - 5 y = 9
================
I suspected that.
Thx for letting us know.
---
PS don't put spaces ...
---------
3x + 5y = 2
9x - 5y = 9
--------------------- Add if you want to solve for x & y.